OpenStudy (anonymous):

What are the coordinates of the vertex for f(x) = 2x2 + 4x + 9? a. (−2, 4) b. (−1, 7) c. (1, 15) d. (2, 9)

OpenStudy (triciaal):

vertex when the slope = 0 can you differentiate the function?

OpenStudy (triciaal):

4x + 4 = 0 solve for x x = -1 what is f(-1)? this is the y value so choice b

OpenStudy (triciaal):

another approach x = -b/2a = -4/2*2 = -4/4 = -1 y = f(-1) = 2(-1)^2 + 4(-1) + 9 = 2 - 4 + 9 =11-4 = 7 choice b